OpenStudy (anonymous):

What is 2 to the power of 4 over 3 equal to?

OpenStudy (anonymous):

2x2x2x2=16 so your answer is 16/3

OpenStudy (anonymous):

so \[^3 \sqrt{16}\]?

OpenStudy (anonymous):

@Poetrylover

OpenStudy (anonymous):

no 16 over three.

OpenStudy (anonymous):

i have multiple choice of \[^3\sqrt{8} \] \[^2\sqrt{8}\] \[^3\sqrt{16}\] and \[^2\sqrt{16}\]

OpenStudy (anonymous):

then i guess 3square root.

OpenStudy (anonymous):

of 16?